# Enhancing Adversarial Robustness of Vision Language Models via Adversarial Mixture Prompt Tuning

### v1

@@ 4.3 Conditional Prompt Weight Router
Algorithm 1 Training Process of AMPT
1:The train dataset $\mathcal{D}$, clean examples $x$ and adversarial examples $\tilde{x}$, ground truth $y$, text encoder $F_{\theta_{t}}$ and image encoder $F_{\theta_{v}}$, adversarial mixture prompts with random initialization $T_{m}=\{t^{1},t^{2},\cdots,t^{K}\}$, total class number $N$, condition prompt weight router $F_{\theta_{w}}$ with random initialization parameter $\theta_{w}$, the max training epochs $max$-$epoch$.
2:for 00 to $max$-$epoch$ do
3:     for $Every~{}minibatch(x,t,y)~{}in~{}\mathcal{D}$ do
4:          $\tilde{x}=\mathop{argmax}\limits_{||\tilde{x}-x||\leq\epsilon}\mathcal{L}_{AMPT}(\tilde{x},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})$.   //  Get adversarial examples $\tilde{x}$.
5:          $\{z_{t,1},\cdots,z_{t,K}\}=\{F_{\theta_{t}}(t^{1}),\cdots,F_{\theta_{t}}(t^{K})\}$.   //  Get each feature of adversarial prompt.
6:         for $each~{}x_{i}~{}in~{}x$ do
7:              $\tilde{z}_{v}^{i}=F_{\theta_{v}}(\tilde{x})$.   //  Get the image feature of adversarial examples $\tilde{x}_{i}$.
8:               $\tilde{w}^{i}=softmax(F_{\theta_{w}}(\tilde{z}_{v}^{i}))$.   //  Obtain the text prompt weights $\tilde{w}^{i}$ based on $\tilde{x}_{i}$.
9:               $\{z_{t}^{1,i},\cdots,z_{t}^{N,i}\}=\{\sum_{k}^{K}\tilde{w}_{k}^{i}z_{t,k}^{1},\cdots,\sum_{k}^{K}\tilde{w}_{k}^{i}z_{t,k}^{N}\}$.   //  Get mixture text features for $\tilde{x}_{i}$.
10:         end for
11:          $\theta_{w}=\theta_{w}-\eta\cdot\nabla_{\theta_{w}}\mathcal{L}_{AMPT}(\tilde{x}_{i},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})$.   //  Update prompt weight router.
12:          $T_{m}=T_{m}-\eta\cdot\nabla_{T_{m}}\mathcal{L}_{AMPT}(\tilde{x}_{i},T_{m},y;F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})$.   //  Update mixture prompts.
13:     end for
14:end for
@@
where $F_{\theta_{w}}$ denotes the conditional prompt weight router, the $z_{v}^{i}$ denotes the feature generated by image encoder of image $x_{i}$. The softmax operation can keep the sum of the weights as 1. The $\tau_{w}$ is applied to control the adjustment strength of the generated weight, while a smaller $\tau_{w}$ denotes stronger adjustment strength, and a larger $\tau_{w}$ denotes weaker adjustment strength, when $\tau_{w}$ approaches infinity, it will degenerate into $1/K$. With the assistance of an adaptive weight router, we can finally more generalizable and representative mixture text feature based on the image features, to further improve the adaptability to defend against different adversarial examples. Meanwhile, we also provide the Theorem 1 about our conditional prompt weight router.
@@
$\mathbb{E}(\sum_{k}^{K}\tilde{w}_{k}\mathcal{R}(\tilde{x},t^{k},y))<\mathbb{E}(\frac{1}{K}\sum_{k}^{K}\mathcal{R}(\tilde{x},t^{k},y)).$ (11)
The proof of Theorem 1 can be viewed in Appendix A.1. Theorem 1 shows that conditional prompt weights can bring the smaller error expectation of the adversarial examples compared with the average error expectation of the adversarial examples, which further demonstrates the necessity and effectiveness of our conditional prompt weight router.
Then the entire process of our AMPT can be viewed in Figure 1, and the optimization loss function $\mathcal{L}_{AMPT}$ can be defined as follows:
$\mathcal{L}_{AMPT}(\tilde{x}_{i},T_{m},y_{i};F_{\theta_{v}},F_{\theta_{t}},F_{\theta_{w}})=-\sum_{j}^{N}y_{ij}log\frac{exp(cos(\tilde{z}_{v}^{i},z_{t}^{j,i}))}{\sum_{j}^{N}exp(cos(\tilde{z}_{v}^{i},z_{t}^{j,i}))},$ (12)
@@ 5.1 Experimental Setting
Datasets. Following [38, 17], we mainly conduct our experiments on 11 high-resolution vision datasets: ImageNet [10], Caltech101 [11], OxfordPets [26], StanfordCars [16], Flowers102 [25], Food101 [3], FGVCAircraft [23], SUN397 [34], DTD [7], EuroSAT [12], and UCF101 [30].
@@
Models. Following the setting in [17], we apply ViT-B/32 as our default selected backbone of image encoder, and select the model trained by a strong AT method TeCoA [24] as our default optimized weight. It should be mentioned that our method can easily transform into other CLIP-liked structures.
@@
Baselines. Because our AMPT is a text prompt tuning method, we mainly compare our method with some similar state-of-the-art methods: Hand Engineered Prompts (HEP), AdvPT [36] and APT [17]. Similar to CoOp [38], we select two versions for AMPT and APT: unified context (UC) and class-specific context (CSC) following APT [17], while AdvPT only contain the UC setting. We change the setting of AdvPT into the setting of APT [17] for the sake of fair comparison. Since AMPT can be well combined with bimodal prompt tuning, e.g., FAP [39] and APD [21], we do not conduct additional comparisons about them.
@@
Evaluation Metric. Here we mainly select two adversarial attacks, PGD attack [22] and AutoAttack [9]. If without additional claim, we set the maximum perturbation $\epsilon$ of adversarial attacks to 4/255. For the PGD attack, we apply 100 iterations with a step $\epsilon/4$ following [17]. Meanwhile, we employ an esemble attack, AutoAttack (AA) [9], which consists of four different attack methods: Auto-PGD (APGD), the Difference of Logits Ratio (DLR) attack, FAB-Attack [8], and the black-box Square Attack [1]. All the methods are evaluated on the entire test test if without additional instruction. For the evaluation of ImageNet against Autoattack, we select the 5000 test set to reduce the calculation overhead following [17], while conducting the AutoAttack on the entire test set is too expensive.
Table 1: Robustness performance(%) under all data training in the Unified Context (UC) setting.
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.84 77.44 61.49 30.37 10.33 7.02 27.13 31.98 21.70 20.31 36.16 33.07 PGD 10.27 44.02 14.28 8.73 0.92 0.48 11.17 5.86 3.19 9.25 6.24 10.40 HEP AA 7.24 39.92 11.01 6.41 0.62 0.06 20.51 3.94 1.76 8.21 4.84 9.50 Clean 44.60 88.88 75.58 81.49 41.47 21.66 53.78 50.34 45.42 79.58 63.44 58.75 PGD 9.05 56.95 12.97 28.46 2.82 2.04 20.27 6.80 5.79 10.37 12.79 15.30 AdvPT-UC [36] AA 7.02 55.13 11.07 24.73 1.62 1.26 18.79 5.50 4.06 9.22 10.73 13.56 Clean 41.48 88.32 72.58 80.88 37.42 20.49 52.19 47.29 35.32 68.67 59.00 54.88 PGD 12.57 63.65 24.56 44.90 8.93 7.05 26.24 13.15 13.11 24.51 21.89 23.69 APT-UC [17] AA 8.16 61.01 16.43 38.61 3.92 3.33 22.40 8.06 7.32 29.79 16.39 19.58 Clean 42.30 87.38 72.72 82.34 45.17 20.58 53.43 51.48 38.98 68.19 60.27 56.62 PGD 12.62 65.03 24.78 46.81 12.16 7.56 28.49 13.92 13.34 37.70 21.99 25.85 MAPT-UC (ours) AA 8.18 62.39 16.57 41.21 6.01 3.21 25.41 9.91 8.03 34.56 17.29 21.16
Table 2: Robustness performance(%) under 16-shot training in the Unified Context (UC) setting.
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.84 77.44 61.49 30.37 10.33 7.02 27.13 31.98 21.70 20.31 36.16 33.07 PGD 10.27 44.02 14.28 8.73 0.92 0.48 11.17 5.86 3.19 9.25 6.24 10.40 HEP AA 7.24 39.92 11.01 6.41 0.62 0.06 20.51 3.94 1.76 8.21 4.84 9.50 Clean 43.09 87.58 73.29 74.46 37.07 19.92 46.45 47.28 36.05 61.40 56.01 52.96 PGD 8.72 50.67 12.84 20.99 2.69 2.07 16.13 6.45 4.31 9.07 10.52 13.13 AdvPT-UC [36] AA 6.72 49.53 10.47 16.89 1.69 0.96 14.54 5.17 2.99 7.36 9.09 11.40 Clean 41.12 86.29 67.29 76.41 31.6 20.31 45.86 44.92 30.39 64.33 53.16 51.06 PGD 12.27 56.75 19.98 37.52 7.70 6.15 21.51 10.94 7.90 25.54 16.55 20.26 APT-UC [17] AA 7.88 53.43 13.46 32.20 3.37 2.64 18.91 6.88 4.17 16.68 12.74 15.67 Clean 41.20 87.14 70.62 76.53 38.54 19.29 47.75 47.32 30.73 57.06 54.09 51.84 PGD 12.38 57.69 20.69 37.79 9.29 6.63 22.99 11.08 8.46 24.72 18.61 20.94 MAPT-UC (ours) AA 7.84 55.05 13.27 32.28 5.07 2.52 20.04 7.75 4.60 19.35 14.14 16.54
@@ 5.2 Robustness Performance
We conduct a benchmark evaluation of our AMPT and baseline approaches. Table 1 and Table 2 present the performance of various text prompt methods in 11 datasets in both full-data and 16-shot training settings. Based on the results, AMPT not only demonstrates strong adversarial robustness under different attacks but also maintains competitive accuracy.
Specifically compared to AdvPT and APT, our approach consistently delivers superior robustness and stronger data efficiency. Under full-data training, AMPT-UC improves robustness over Baseline by an average of 2.16% and 1.58% (PGD/AA). Under the 16-shot training, AMPT-UC surpasses the Baseline by an average of 0.68% / 0.87% (PGD/AA), highlighting that the AMPT enhances APT’s performance under different types of attacks.
We also compare the two variants of APT-CSC and AMPT-CSC in Table 3. AMPT-CSC outperforms APT-CSC by an average of 0.27% / 0.23% (PGD/AA). This is likely because the CSC variant contains more parameters and therefore has a higher capacity for both APT-CSC and AMPT-CSC, and the insufficient generalization of APT-CSC will be somewhat alleviated, but AMPT-CSC still achieves better robustness, showing the effectiveness.
Table 3: Robustness performance(%) under all data training in Class-Specific Context (CSC) setting.
@@
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.74 89.49 72.74 87.82 49.04 31.44 55.56 51.00 42.62 69.77 64.47 59.43 PGD 11.85 65.44 24.56 55.30 12.92 9.00 27.54 13.74 14.71 38.32 24.16 27.05 APT-CSC[17] AA 8.26 63.41 17.77 50.14 6.75 5.01 24.35 9.57 9.04 33.40 19.69 22.49 Clean 38.85 88.56 72.58 86.40 45.28 27.84 55.50 51.83 42.82 69.79 62.86 58.39 PGD 11.83 66.82 25.73 53.92 12.39 9.30 29.20 14.01 14.71 38.77 23.82 27.32 MAPT-CSC(ours) AA 7.94 64.67 18.12 49.05 6.50 4.86 25.83 9.88 9.13 34.65 19.30 22.72
5.3 Generalization across Different Datasets
Table 4: Comparsion of Robust Generalization (%) between APT-UC and AMPT-UC cross 9 different datasets.
Metric APT-UC AMPT-UC(ours)
Clean 14.75 18.39
PGD 3.39 4.89
AA 2.66 4.14
@@ 5.3 Generalization across Different Datasets
Meanwhile, we test the robust generalization across different datasets. We apply the APT-UC and AMPT-UC trained on Caltech101 with all-data training setting as source models and evaluate the adversarial robustness in different datasets, including OxfordPets, OxfordFlowers, StanfordCars, FGVCAircraft, Sun397, DTD, Food101, EruoSAT, and UCF101, and the average results are reported in Table 4. From the result, we find that our AMPT-UC has better robust generalization compared with APT-UC. Specifically, Our AMPT-UC has a 1.5% and 1.78% robustness improvement against PGD and AA attacks compared with APT-UC, showing that our AMPT has better performance in dealing with diverse adversarial examples even in unseen classes. The detailed results can be viewed in Appendix A.2.

### v2

@@ 4.3 Conditional Prompt Weight Router
where $F_{\theta_{w}}$ denotes the conditional prompt weight router, the $z_{v}^{i}$ denotes the feature generated by image encoder of image $x_{i}$. The softmax operation can keep the sum of the weights as 1. The $\tau_{w}$ is applied to control the adjustment strength of the generated weight, while a smaller $\tau_{w}$ denotes stronger adjustment strength, and a larger $\tau_{w}$ denotes weaker adjustment strength, when $\tau_{w}$ approaches infinity, it will degenerate into $1/K$. With the assistance of an adaptive weight router, we can finally obtain a more generalizable and representative mixture text feature based on the image features, to further improve the adaptability to defend against different adversarial examples. Meanwhile, we also provide the Theorem 1 about our conditional prompt weight router.
Algorithm 1 Training Process of MoAPT
1:The train dataset $\mathcal{D}$, clean examples $x$ and adversarial examples $\tilde{x}$, ground truth $y$, text encoder $F_{\theta_{t}}$ and image encoder $F_{\theta_{v}}$, adversarial mixture prompts with random initialization $T_{m}=\{t^{1},t^{2},\cdots,t^{K}\}$, total class number NN, condition prompt weight router $F_{\theta_{w}}$ with parameter $\theta_{w}$, the max training epochs maxmax-epochepoch, the router temperature $\tau_{w}$
2:for 0 to maxmax-epochepoch do
3:  for $Every~minibatch(x,t,y)~in~\mathcal{D}$ do
[…]
5:    $\{z_{t,1},\cdots,z_{t,K}\}=\{F_{\theta_{t}}(t^{1}),\cdots,F_{\theta_{t}}(t^{K})\}$.
6:   for $each~x_{i}~in~x$ do
[…]
8:      $\tilde{w}^{i}=softmax(F_{\theta_{w}}(\tilde{z}_{v}^{i})/\tau_{w})$.
9:      $z_{t}^{n,i}=\sum_{k}^{K}\tilde{w}_{k}^{i}z_{t,k}^{n}$.
10:   end for
11:    $\theta_{w}=\theta_{w}-\eta\cdot\nabla_{\theta_{w}}\mathcal{L}_{MoAPT}$.
12:    $T_{m}=T_{m}-\eta\cdot\nabla_{T_{m}}\mathcal{L}_{MoAPT}$.
13:  end for
14:end for
@@
$\mathbb{E}(\sum_{k}^{K}\tilde{w}_{k}\mathcal{R}(\tilde{x},t^{k},y))\leq\mathbb{E}(\frac{1}{K}\sum_{k}^{K}\mathcal{R}(\tilde{x},t^{k},y)),$ (11)
when there exists at least one pair $(i,j)$ exists $i\neq j$, such that $\mathcal{R}(\tilde{x},t^{i},y))<\mathcal{R}(\tilde{x},t^{j},y))$, the strict inequality holds.
The proof of Theorem 1 can be viewed in Appendix 1.1. Theorem 1 shows that conditional prompt weights can bring the smaller error expectation of the adversarial examples compared with the average error expectation of the adversarial examples, which further demonstrates the necessity and effectiveness of our conditional prompt weight router.
Then the entire process of our MoAPT can be viewed in Figure 1, and the optimization loss function $\mathcal{L}_{MoAPT}$ can be defined as follows:
$\mathcal{L}_{MoAPT}=-\sum_{n}^{N}y_{in}log\frac{exp(cos(\tilde{z}_{v}^{i},z_{t}^{n,i}))}{\sum_{n}^{N}exp(cos(\tilde{z}_{v}^{i},z_{t}^{n,i}))},$ (12)
@@ 5 Experiments
Table 1: Robustness performance(%) with all data training setting on 11 different datasets under maximum perturbation 4/255.
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.84 77.44 61.49 30.37 10.33 7.02 27.13 31.98 21.70 20.31 36.16 33.07 PGD 10.27 44.02 14.28 8.73 0.92 0.48 11.17 5.86 3.19 9.25 6.24 10.40 HEP AA 7.24 39.92 11.01 6.41 0.62 0.06 9.52 3.94 1.76 8.21 4.84 9.50 Clean 48.84 84.63 62.25 67.19 1.07 0.99 22.05 48.91 39.89 78.89 12.11 42.44 PGD 5.78 51.36 15.51 34.51 0.91 0.99 9.99 17.48 14.15 51.70 4.57 18.81 VPT [24] AA 1.44 11.52 0.05 1.79 0.65 0.00 1.77 0.51 0.49 5.26 0.32 2.16 Clean 52.17 91.03 80.07 86.43 50.21 23.88 60.81 58.35 64.38 89.71 68.25 65.94 PGD 7.39 54.27 12.78 27.81 2.11 1.32 20.74 6.74 6.67 22.66 14.53 16.09 FAP [39] AA 0.89 11.88 1.17 2.67 0.27 0.39 8.09 0.74 0.94 19.23 1.71 4.36 Clean 44.60 88.88 75.58 81.49 41.47 21.66 53.78 50.34 45.42 79.58 63.44 58.75 PGD 9.05 56.95 12.97 28.46 2.82 2.04 20.27 6.80 5.79 10.37 12.79 15.30 AdvPT [36] AA 7.02 55.13 11.07 24.73 1.62 1.26 18.79 5.50 4.06 9.22 10.73 13.56 Clean 41.48 88.32 72.58 80.88 37.42 20.49 52.19 47.29 35.32 68.67 59.00 54.88 PGD 12.57 63.65 24.56 44.90 8.93 7.05 26.24 13.15 13.11 24.51 21.89 23.69 APT [17] AA 8.16 61.01 16.43 38.61 3.92 3.33 22.40 8.06 7.32 29.79 16.39 19.58 Clean 42.30 87.38 72.72 82.34 45.17 20.58 53.43 51.48 38.98 68.19 60.27 56.62 PGD 12.62 65.03 24.78 46.81 12.16 7.56 28.49 13.92 13.34 37.70 21.99 25.85 MoAPT (ours) AA 8.18 62.39 16.57 41.21 6.01 3.21 25.41 9.91 8.03 34.56 17.29 21.16
@@ 5.1 Experimental Setting
Datasets. Following [17], we conduct our experiments on 11 high-resolution vision datasets: ImageNet [10], Caltech101 [11], OxfordPets [26], StanfordCars [16], Flowers102 [25], Food101 [3], FGVCAircraft [23], SUN397 [34], DTD [7], EuroSAT [12], and UCF101 [30]. The 11 datasets were selected to establish a comprehensive benchmark, covering a wide range of vision tasks including generic object classification, scene recognition, action classification, fine-grained recognition, texture recognition, and satellite imagery analysis. They were split into training and test sets following the protocol of  [38].
@@
Models. Following the setting in [17], we apply ViT-B/32 as our default selected backbone of image encoder, and select the model trained by a strong AT method TeCoA [24] as our default optimized weight.
@@
Baselines. Because our MoAPT is a text prompt tuning method, we mainly compare our method with some similar state-of-the-art methods: Hand Engineered Prompts (following the setting in [17], see Appendix 1.3 for details), VPT [24], AdvPT [36], APT [17], FAP [39], where VPT is a visual prompt tuning method, AdvPT and APT are text prompt tuning method, FAP is the bi-modal tuning method. Here we apply the HEP following the setting in [17]. Meanwhile, we change the setting of AdvPT into the setting of APT [17] for the sake of fair comparison. To ensure fairness, we apply the same backbone to further enhance the robustness for all the baselines.
@@
Evaluation Metric. Following the setting in [17], we select two adversarial attacks, PGD attack [22] and AutoAttack [9]. If without additional claim, we set the maximum perturbation $\epsilon$ of adversarial attacks to 4/255. For the PGD attack, we apply 100 iterations with a step $\epsilon/4$ following [17]. Meanwhile, we employ an ensemble attack, AutoAttack (AA) [9], which consists of four different attack methods: Auto-PGD (APGD), the Difference of Logits Ratio (DLR) attack, FAB-Attack [8], and the black-box Square Attack [1]. All the methods are evaluated on the entire test test if without additional instruction. For the evaluation of ImageNet against Autoattack, we select the 5000 test set to reduce the calculation overhead following [17], while conducting the AutoAttack on the entire test set is too expensive.
@@ 5.2 Robustness Performance
Table 2: Robustness performance(%) with 16-shot training setting on 11 different datasets under maximum perturbation 4/255.
@@
Methods Metric ImageNet Caltech101 OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 39.84 77.44 61.49 30.37 10.33 7.02 27.13 31.98 21.70 20.31 36.16 33.07 PGD 10.27 44.02 14.28 8.73 0.92 0.48 11.17 5.86 3.19 9.25 6.24 10.40 HEP AA 7.24 39.92 11.01 6.41 0.62 0.06 9.52 3.94 1.76 8.21 4.84 9.50 Clean 34.84 76.92 3.38 41.49 3.38 1.05 11.64 44.02 1.16 6.70 2.11 20.61 PGD 3.13 28.28 0.25 13.85 0.25 0.93 0.71 13.39 0.08 0.00 0.13 5.55 VPT [24] AA 0.71 0.30 0.14 0.20 0.14 0.00 0.24 0.30 0.09 0.10 0.19 0.22 Clean 50.34 89.85 76.09 76.24 43.68 19.44 50.29 56.24 55.39 64.67 63.75 58.73 PGD 6.92 49.77 11.28 17.45 1.67 1.32 16.72 2.44 4.21 13.48 10.46 12.34 FAP [39] AA 0.51 8.92 1.74 1.46 0.18 0.21 6.85 1.40 0.66 11.17 0.97 3.10 Clean 43.09 87.58 73.29 74.46 37.07 19.92 46.45 47.28 36.05 61.40 56.01 52.96 PGD 8.72 50.67 12.84 20.99 2.69 2.07 16.13 6.45 4.31 9.07 10.52 13.13 AdvPT [36] AA 6.72 49.53 10.47 16.89 1.69 0.96 14.54 5.17 2.99 7.36 9.09 11.40 Clean 41.12 86.29 67.29 76.41 31.6 20.31 45.86 44.92 30.39 64.33 53.16 51.06 PGD 12.27 56.75 19.98 37.52 7.70 6.15 21.51 10.94 7.90 25.54 16.55 20.26 APT [17] AA 7.88 53.43 13.46 32.20 3.37 2.64 18.91 6.88 4.17 16.68 12.74 15.67 Clean 41.20 87.14 69.58 77.63 38.54 19.29 47.75 47.32 30.73 57.06 54.09 51.84 PGD 12.38 57.69 21.29 39.14 9.29 6.63 22.99 11.08 8.46 28.96 18.61 21.50 MoAPT (ours) AA 7.84 55.05 14.39 34.02 5.07 2.52 20.04 7.75 4.60 19.35 14.14 16.80
We conduct a benchmark evaluation of our MoAPT and baseline approaches. Table 1 and Table 2 present the performance of various prompt methods in 11 datasets in both full-data and 16-shot training settings. Based on the results, MoAPT improves robustness by an average of 8.99% and 11.33% (PGD/AA) in the all-shot setting, and by 9.17% and 8.82% (PGD/AA) in the 16-shot setting. Furthermore, MoAPT achieves an average accuracy improvement of 5.60% and 8.55% (under all /16 shots training settings). It demonstrates strong adversarial robustness across various attacks while maintaining competitive accuracy.
Specifically, MoAPT consistently outperforms AdvPT and the best baseline APT in robustness and data efficiency. Under full-data training, MoAPT improves robustness over the best baseline by 2.16% and 1.58% (PGD/AA). In the 16-shot setting, MoAPT surpasses the baseline by 1.24% and 1.13% (PGD/AA), demonstrating its enhancement of APT’s performance under various attacks. In contrast to MoAPT, VPT and FAP perform poorly in evaluating AA attacks, likely due to their lack of the generalization ability to unseen attacks, as seen in MoAPT.
Table 3: Out-of-Distribution Robustness (%) between APT and MoAPT cross 9 different datasets based on Caltech101 adversarial prompts.
Method Metric OxfordPets Flowers102 Cars FGVC DTD SUN397 Food101 EuroSAT UCF101 Average Clean 29.95 14.01 8.10 1.83 16.84 14.19 15.52 12.81 19.51 14.75 PGD 10.06 3.17 0.87 0.36 7.74 2.25 1.43 1.14 3.49 3.39 APT AA 8.67 2.15 0.39 0.33 6.74 1.63 0.77 0.58 2.72 2.66 Clean 45.08 16.00 13.01 2.34 16.02 20.17 15.3 11.43 26.20 18.39 PGD 9.10 4.02 1.02 0.81 7.68 4.10 2.24 10.67 4.34 4.89 MoAPT(ours) AA 7.14 2.84 0.53 0.75 6.97 4.02 1.31 10.57 3.09 4.14
@@
Meanwhile, we test the out-of-distribution robustness across different datasets. We apply the APT and MoAPT trained on Caltech101 with all-data training setting as source models and evaluate the adversarial robustness in different datasets, including OxfordPets, OxfordFlowers, StanfordCars, FGVCAircraft, Sun397, DTD, Food101, EuroSAT, and UCF101, and the average results are reported in Table 3. From the result, we find that MoAPT has better robust generalization compared with APT. Specifically, MoAPT has a 1.5% and 1.48% robustness improvement against PGD and AA attacks compared with APT, showing that MoAPT has better performance in dealing with diverse adversarial examples even in unseen classes. MoAPT also outperforms in different backbone models, see Appendix 1.4 for details.
5.3 Trade-off between Accuracy and Robustness
As shown in Fig. 3, we compare the performance improvement per dataset of our adversarially-trained prompt over the standard-trained prompt for unified context. Most adversarially trained vision models tend to improve robustness at the expense of accuracy, and adversarially trained prompts also exhibit this trade-off, which is expected. More importantly, we observe that for most datasets, the gain in robustness outweighs the drop in accuracy. Specifically, MoAPT improves adversarial robustness by an average of +7.60%, while incurring only a modest drop of -3.03% in accuracy. For instance, on OxfordPets, robustness increases significantly by +8.61%, with a slight gain of +0.23% in accuracy. These results suggest that our method achieves a relatively favorable trade-off between accuracy and robustness.
Figure 3: Trade-off between Accuracy and Robustness ($M=16$).